# Generating Directed Graphs with Dual Attention and Asymmetric Encoding

### v2

@@ 1 Introduction
Figure 1: (a) The learnable space increases drastically with the number of nodes for digraphs compared to undirected graphs (Harary & Palmer, 1973), highlighting challenges in extending graph generative models to directed structures. (b) Overview of our generative model for directed graphs. During training, the dual attention transformer (denoising network), parameterized by $\theta$, is enhanced with asymmetric positional encoding, learning to reverse predictions using cross-entropy loss. During inference, we compute the rate matrix $R_{t}^{\theta}(G_{t},G_{t+\Delta t})$, which governs the evolution of the generative process over finite intervals $\Delta t$, based on the model prediction $p^{\theta}(\cdot\mid G_{t})$.
@@
To address the first challenge, we introduce Directo  [^1] Code available at: https://anonymous.4open.science/r/DirectoAnonymous, the first iterative refinement-based generative model for directed graphs. Our approach brings together three core components of expressive graph generative modeling: (i) an expressive neural architecture, (ii) a robust generative framework, and (iii) informative input features. For the architecture, we employ a dual attention mechanism that performs cross-attention between edge features and their reversed counterparts, capturing both source-to-target and target-to-source information flow. All this is integrated within a robust discrete-state flow matching framework that enables efficient generation with state-of-the-art performance and is also readily extendable to conditional generation tasks through classifier-free guidance (Ho & Salimans, 2022). Finally, the network’s input features include positional encodings designed for asymmetric adjacency matrices, which we show outperform directionality-agnostic alternatives. Figure 1(b) gives an overview of our method and its components.
@@ 2 Background: Discrete Flow Matching for Graph Generation
We denote by $G=\left(x^{(1:n:N)},\,e^{(1:i\neq j:N)}\right)$ directed graphs with $N$ nodes. The set of nodes is denoted by $\{x^{(n)}\}_{n=1}^{N}$, and the set of directed edges by $\{e^{(i,j)}\}_{1\leq i\neq j\leq N}$. Both nodes and edges are categorical variables, where $x^{(n)}\in\{1,\ldots,X\}$ and $e^{(i,j)}\in\{1,\ldots,E\}$. Inspired by standard practice in iterative refinement for undirected graphs (Vignac et al., 2023a; Xu et al., 2024; Siraudin et al., 2024; Qin et al., 2025a), we assume that every ordered pair of distinct nodes corresponds to a directed edge in one of several possible classes, including a class representing an absent edge. Finally, we use a subscript $t$ to denote variables at time $t$, e.g., $x_{t}^{(n)}$ or $e_{t}^{(i,j)}$.
@@
The noising process runs from $t=1$ to $t=0$, progressively corrupting the original input graphs through a linear interpolation between the data distribution and the limit distribution $\bm{p}_{0}=\bm{p}_{\text{noise}}$ (as detailed in Figure 1(b)). This interpolation is applied independently to each variable (Gat et al., 2024; Campbell et al., 2024), corresponding to each node and edge in the case of graphs. For example, for each node $x_{1}^{(n)}$, its noisy distribution at step $t$ is defined as:
@@
where $\bm{R}_{t}$ is the CTMC rate matrix (Campbell et al., 2024). In practice, this update rule is approximated in two ways: first, by discretizing time with a finite step size $\Delta t$, in an Euler method step; and second, by approximating the ground-truth rate matrix with an estimate $\bm{R}^{\theta}_{t}$ computed from the predictions of a denoising neural network $\bm{p}^{\theta}_{1\mid t}(\cdot\mid G_{t})$ parametrized by $\theta$. This network is trained to predict the clean node $\bm{p}_{1\mid t}^{\theta,(n)}$ and edge $\bm{p}_{1\mid t}^{\theta,(i,j)}$ distributions given a noisy graph, aggregated as:
$\bm{p}^{\theta}_{1|t}(\cdot|G_{t})=\left(\left(\bm{p}_{1\mid t}^{\theta,(n)}(x_{1}^{(n)}\mid G_{t})\right)_{1\leq n\leq N},\left(\bm{p}_{1\mid t}^{\theta,(i,j)}(e_{1}^{(i,j)}\mid G_{t})\right)_{1\leq i\neq j\leq N}\right).$ (3)
@@ 3 Generating DiGraphs with Asymmetric Encoding and Dual Attention
In this section, we introduce Directo, the first flow-based digraph generative model. We begin by outlining the overall generative framework, followed by the two directionality-aware components that form the core of Directo: asymmetric positional encoding and dual attention.
@@ 3.1 Directed Graph Generation via Discrete Flow Matching
3.2 Asymetric positional encoding
@@
We hypothesize that incorporating directionality-aware positional encodings improves digraph generative performance. To test this, we design an ablation study evaluating which of the described encodings provide the most informative structural signals, while also considering their computational cost, to identify the best tradeoff between performance and efficiency.
@@ 3.3 Graph transformer with dual attention
Graph Transformers (Dwivedi & Bresson, 2021; Rampášek et al., 2022) effectively model node interactions by leveraging the adjacency matrix to encode structural information. However, in directed graphs, it is crucial for the model to distinguish the different semantics carried by incoming and outgoing edges, which makes the standard graph transformer insufficiently expressive. Generating digraphs thus requires models that (i) capture bidirectional information flow to learn how source and target nodes influence each other distinctly, and (ii) integrate structural signals across nodes, edges, and the global graph to preserve relational dependencies and ensure coherent outputs.
@@
To model directional dependencies more effectively, we use a cross-attention mechanism between source-to-target (outgoing) edge features $\bm{E}_{\text{ST}}$ and their reversed target-to-source (incoming) counterparts $\bm{E}_{\text{TS}}$. By explicitly attending to both directions, our architecture, which builds on ideas from Wang et al. (2024), processes information bidirectionally and better captures directed relationships, improving expressiveness and accuracy. Concretely, we compute two directional attention maps between role-specific node projections:
@@
where $h$ is the hidden dimension, $\bm{V}_{\text{S}},\bm{V}_{\text{T}}$ are the value vectors for source and target nodes, and the concatenation operation is performed along the node dimensions. These unified weights are then used to compute a convex (i.e., softmax-weighted) linear combination of the value vectors:
@@ 4.1 Datasets
A relevant application of DAGs is Neural Architecture Search, which aims to generate novel instances to achieve an automated design of efficient Neural Networks (Elsken et al., 2019). To support this task, the TPU Tiles dataset (Phothilimthana et al., 2023) contains computational graphs extracted from Tensor Processing Unit workloads. We split into 5,040 training, 630 validation, and 631 test graphs following Li et al. (2025).
@@ 4.2 Metrics
To evaluate proximity to the original distribution, we compute the Maximum Mean Discrepancy (MMD) (Martinkus et al., 2022) between generated and test graphs across several statistics: outgoing and incoming degree distributions, directed clustering coefficients, spectral, and wavelet-based features. Spectral features are derived from the directed Laplacian (Chung, 2005), simplifying computations compared to the Magnetic Laplacian. We normalize results as the ratio from the generated samples MMD to that of the training data and then average the ratios.
@@ 5 Experiments
In the general directed setting, we compare Directo against performing Maximum Likelihood Estimation (MLE) baseline of the node count, node types, and edge types, and sampling from the resulting distribution to generate new graph instances. For DAG settings, we also include the only two publicly available baselines: D-VAE (Zhang et al., 2019) and LayerDAG (Li et al., 2025). Notably, the architectural improvements in Sections 3.2 and 3.3 are agnostic to the underlying iterative refinement framework. To demonstrate versatility and enable wider experimentation, we extend Directo to a discrete diffusion (Vignac et al., 2023a) backbone (Directo-DD), described in Section I.3. We also include DeFoG (Qin et al., 2025a) and DiGress (Vignac et al., 2023a), discrete flow matching and diffusion methods in the undirected setting, respectively, adapting both to the directed case by removing the edge symmetrization operations. For each experiment, we highlight the best result and second-best method, with further experimental details in Appendix G.
@@ 5.1 Generative Performance Evaluation
As groundwork for directed graph generation, we prioritize strong unconditional generation, which we consider a key prerequisite for effective digraph modeling. At the same time, steering generation with graph-level properties is important for many downstream tasks. To illustrate the flexibility of our framework, we show that Directo can be extended to conditional generation without major architectural changes via classifier-free guidance (Ho & Salimans, 2022; Nisonoff et al., 2025). On the TPU Tiles dataset, we condition on the execution time of the computational graph, aiming at low execution times. Directo performs competitively with specialized autoregressive models like GraphRNN and LayerDAG (Appendix H.7), even surpassing the OneShotDAG variant of Li et al. (2025).
@@ 5.2 Ablation study
width= Dataset Model Ratio $\downarrow$ V.U.N. $\uparrow$ ER-DAG RRWP - Double 4.9 $\pm$ 0.2 72.0 $\pm$ 9.8 RRWP - Dual 1.7 $\pm$ 0.1 94.0 $\pm$ 1.0 MagLap - Double 4.3 $\pm$ 0.7 80.0 $\pm$ 6.3 MagLap - Dual 1.3 $\pm$ 0.2 91.0 $\pm$ 2.5 SBM RRWP - Double 27.1 $\pm$ 2.3 0.0 $\pm$ 0.0 RRWP - Dual 1.4 $\pm$ 0.2 87.0 $\pm$ 5.1 MagLap - Double 24.8 $\pm$ 4.6 6.0 $\pm$ 4.9 MagLap - Dual 1.9 $\pm$ 0.3 77.0 $\pm$ 7.6
@@ 6 Related work
Initial methods include auto-regressive models (You et al., 2018; Liao et al., 2019), which grow the graph progressively by inserting nodes and edges, but require node ordering for feasibility. On the other hand, one-shot graph generative models such as VAEs (Kipf & Welling, 2016; Simonovsky & Komodakis, 2018; Jin et al., 2018; Ma & Zhang, 2021; Vignac & Frossard, 2022), GANs (De Cao & Kipf, 2018; Krawczuk et al., 2021; Martinkus et al., 2022), and normalizing flows (Kaushalya et al., 2019; Wehenkel & Louppe, 2021; Lippe & Gavves, 2021; Luo et al., 2021b) can directly generate full graphs in a single forward pass, thereby eliminating the need for predefined node ordering. Graph diffusion models also belong to this class, with the initial approaches consisting of adaptations of continuous state-space discrete-time diffusion frameworks (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song & Ermon, 2019) to the graph setting. These include works such as EDP-GNN (Niu et al., 2020), DGSM (Luo et al., 2021a), and GeoDiff (Xu et al., 2022). Later, the discrete state-space diffusion framework (Austin et al., 2021) was adopted, with DiGress (Vignac et al., 2023a) and MCD (Haefeli et al., 2022) operating in discrete time. On the other hand, methods that operate in continuous time (Song et al., 2021; Campbell et al., 2022) were introduced, including GDSS (Jo et al., 2022) and GruM (Jo et al., 2024) in continuous state-spaces and DisCo (Xu et al., 2024) and Cometh (Siraudin et al., 2024) for discrete state-spaces. Recent approaches combine the sequential modeling of auto-regressive methods with the global denoising of diffusion (Zhao et al., 2024). Finally, DeFoG (Qin et al., 2025a) recently achieves state-of-the-art performance by leveraging discrete flow matching (Campbell et al., 2024; Gat et al., 2024).
@@ 7 Conclusion and future directions
Despite these promising results, limitations remain. First, scalability can be further improved in handling larger datasets and graph sizes (a direct result of the combinatorial complexity of digraphs). Incorporating scalable strategies, such as sparsity-aware attention (Qin et al., 2025b), hierarchical (Bergmeister et al., 2024; Jang et al., 2024), or latent diffusion (Yang et al., 2024), could help extend Directo at scale. Second, although our current setup includes classifier-free conditional generation, exploring alternative conditional strategies could further improve performance. Third, while Directo can learn structural properties like acyclicity implicitly, enforcing such constraints (e.g., via PRODIGY (Sharma et al., 2024) or ConStruct (Madeira et al., 2024)) could further improve controllability and domain-specific validity. Exploring these directions would enhance the applicability and impact of digraph generative models, particularly on new real-world scenarios. Appendix K includes a detailed discussion of limitations and future work.

### v3

@@ 1 Introduction
Figure 1: (a) The learnable space increases drastically with the number of nodes for digraphs compared to undirected graphs (Harary and Palmer, 1973), highlighting challenges in extending graph generative models to directed structures. (b) Overview of our generative model for directed graphs (with node colors representing different classes). During training, the dual attention transformer (denoising network), parameterized by $\theta$, is enhanced with asymmetric positional encoding, learning to reverse predictions using cross-entropy loss. During inference, we compute the rate matrix $R_{t}^{\theta}(G_{t},G_{t+\Delta t})$, which governs the evolution of the generative process over finite intervals $\Delta t$, based on the model prediction $p^{\theta}(\cdot\mid G_{t})$.
@@
To address the first challenge, we introduce Directo  [^1] Code available at: https://github.com/acarballocastro/DIRECTO, the first iterative refinement-based generative model for directed graphs of small to medium scale (from a few up to 200 nodes on average). Our approach brings together three core components of expressive graph generative modeling: (i) an expressive neural architecture, (ii) a robust generative framework, and (iii) informative input features. For the architecture, we employ a dual attention mechanism that performs cross-attention between edge features and their reversed counterparts, capturing both source-to-target and target-to-source information flow. All this is integrated within a robust discrete-state flow matching framework that enables efficient generation with state-of-the-art performance and more efficiently than diffusion-based alternatives, and which is also readily extendable to conditional generation tasks through classifier-free guidance (Ho and Salimans, 2022). Finally, the network’s input features include positional encodings designed for asymmetric adjacency matrices, which we show outperform directionality-agnostic alternatives. Figure 1(b) gives an overview of our method and its components.
@@ 2 Background: Discrete Flow Matching for Graph Generation
We denote by $G=\left(x^{(1:n:N)},\,e^{(1:i\neq j:N)}\right)$ directed graphs with $N$ nodes. The set of nodes is denoted by $\{x^{(n)}\}_{n=1}^{N}$, and the set of directed edges by $\{e^{(i,j)}\}_{1\leq i\neq j\leq N}$. Both nodes and edges are categorical variables, where $x^{(n)}\in\{1,\ldots,X\}$ and $e^{(i,j)}\in\{1,\ldots,E\}$. Inspired by standard practice in iterative refinement for undirected graphs (Vignac et al., 2023a; Xu et al., 2024; Siraudin et al., 2024; Qin et al., 2025a), we assume that every ordered pair of distinct nodes corresponds to a directed edge. Each edge is treated separately and always belongs to one of several possible categorical classes, including a class representing an absent edge. This formulation allows edges $(i,j)$ and $(j,i)$ to take different classes: both may be absent (no connection between nodes $i$ and $j$), only one may exist, or both may exist, even with different classes. Finally, we use a subscript $t$ to denote variables at time $t$, e.g., $x_{t}^{(n)}$ or $e_{t}^{(i,j)}$.
@@
The noising process runs from $t=1$ to $t=0$, progressively corrupting the original input graphs through a linear interpolation between the data distribution and the limit distribution $\bm{p}_{0}=\bm{p}_{\text{noise}}$ (as detailed in Figure 1(b)). This interpolation is applied independently to each variable following discrete iterative refinement frameworks (Austin et al., 2021b; Campbell et al., 2022; Gat et al., 2024; Campbell et al., 2024), which corresponds to each node and edge in the case of graphs Qin et al. (2025a). For example, for each node $x_{1}^{(n)}$, its noisy distribution at step $t$ is defined as:
@@
where $\bm{R}_{t}$ is the CTMC rate matrix (Campbell et al., 2024). In practice, this update rule is approximated in two ways: first, by discretizing time with a finite step size $\Delta t$, in an Euler method step; and second, by approximating the ground-truth rate matrix with an estimate $\bm{R}^{\theta}_{t}$ computed from the predictions of a denoising neural network $\bm{p}^{\theta}_{1\mid t}(\cdot\mid G_{t})$ parametrized by $\theta$. The denoising neural network, typically a denoising graph transformer, is trained to predict the clean node $\bm{p}_{1\mid t}^{\theta,(n)}$ and edge $\bm{p}_{1\mid t}^{\theta,(i,j)}$ distributions given a noisy graph, aggregated as:
$\bm{p}^{\theta}_{1|t}(\cdot|G_{t})=\left(\left(p_{1\mid t}^{\theta,(n)}(x_{1}^{(n)}\mid G_{t})\right)_{1\leq n\leq N},\left(p_{1\mid t}^{\theta,(i,j)}(e_{1}^{(i,j)}\mid G_{t})\right)_{1\leq i\neq j\leq N}\right).$ (3)
Intuitively, the denoising graph transformer predicts larger probability of the clean graph $\bm{p}_{1\mid t}^{\theta}(\cdot\mid G_{t})$ given the noisy graph $G_{t}$, which in turn is used to compute a higher transition rate in the rate matrix $\bm{R}^{\theta}_{t}$ (see Equation 43 for further clarification). This means that the network predicts the most likely transitions required to denoise the graph and recover the clean target, directly influencing the estimated rate of change.
@@ 3 Generating DiGraphs with Asymmetric Encoding and Dual Attention
Building on the discrete diffusion setting previously described, we now introduce Directo, the first flow-based digraph generative model. We begin by outlining the overall generative framework, followed by a detailed description of the two directionality-aware components that form the core of Directo: the use of directed graph positional encodings and the dual attention mechanism to enhance the edge directionality awareness of the denoising neural network in our framework.
@@ 3.1 Directed Graph Generation via Discrete Flow Matching
3.2 Asymmetric positional encoding
@@
We hypothesize that incorporating directionality-aware positional encodings improves digraph generative performance. To test this, we design an ablation study (see Section 5.2, with further details in Appendix H) evaluating which of the described encodings provide the most informative structural signals, while also considering their computational cost, to identify the best tradeoff between performance and efficiency.
@@ 3.3 Graph transformer with dual attention
Graph Transformers (Dwivedi and Bresson, 2021; Rampášek et al., 2022) effectively model node interactions by leveraging the adjacency matrix to encode structural information. However, in directed graphs, it is crucial for the model to distinguish the different semantics carried by incoming and outgoing edges, which makes the standard graph transformer insufficiently expressive. Generating digraphs thus requires models that (i) capture bidirectional information flow to learn how source and target nodes influence each other distinctly, and (ii) integrate structural signals across nodes, edges, and the global graph to preserve relational dependencies and ensure coherent outputs.
@@
To model directional dependencies more effectively, we use a cross-attention mechanism between source-to-target (outgoing) edge features $\bm{E}_{\text{ST}}$ and their reversed target-to-source (incoming) counterparts $\bm{E}_{\text{TS}}$. By explicitly attending to both directions, our architecture, which builds on ideas from Wang et al. (2024), processes information bidirectionally and better captures directed relationships, improving accuracy and resulting in better generations. Concretely, we compute two directional attention maps between role-specific node projections:
@@
where $h$ is the hidden dimension, $\bm{V}_{\text{S}},\bm{V}_{\text{T}}$ are the value vectors for source and target nodes, and the concatenation operation is performed along the node dimensions. These unified weights are then used to compute a convex (i.e., softmax-weighted) combination of the value vectors:
@@ 4.1 Datasets
A relevant application of DAGs is Neural Architecture Search, which aims to generate novel instances to achieve an automated design of efficient Neural Networks (Elsken et al., 2019). To support this task, the TPU Tiles dataset (Phothilimthana et al., 2023) contains computational graphs extracted from Tensor Processing Unit workloads. We split into 5,040 training, 630 validation, and 631 test graphs following the preprocessing from Li et al. (2025).
@@ 4.2 Metrics
To evaluate proximity to the original distribution, we compute the Maximum Mean Discrepancy (MMD) (Martinkus et al., 2022) between generated and test graphs across several statistics: outgoing and incoming degree distributions, directed clustering coefficients, spectral, and wavelet-based features. Spectral features are derived from the directed Laplacian (Chung, 2005), simplifying computations compared to the Magnetic Laplacian. Following standard practice, we first normalize results as the ratio from the generated samples MMD to that of the training data and then average the ratios for summarization. Values per MMD descriptor are detailed in the extended results in Appendix H..
@@ 5 Experiments
In the general directed setting, we compare Directo against performing Maximum Likelihood Estimation (MLE) baseline of the node count, node types, and edge types, and sampling from the resulting distribution to generate new graph instances. For DAG settings, we also include the only two publicly available baselines: D-VAE (Zhang et al., 2019) and LayerDAG (Li et al., 2025). Notably, the architectural improvements in Sections 3.2 and 3.3 are agnostic to the underlying iterative refinement framework. To demonstrate versatility and enable wider experimentation, we extend Directo to a discrete diffusion (Vignac et al., 2023a) backbone (Directo-DD), described in Section I.3. We also include DeFoG (Qin et al., 2025a) and DiGress (Vignac et al., 2023a), discrete flow matching and diffusion methods in the undirected setting, respectively, adapting both to the directed case by removing the edge symmetrization operations. For each experiment, we highlight the best result and second-best method, with further experimental details in Appendix G, including Table 8 for a fairness comparison of resources and runtime of each of the models.
@@ 5.1 Generative Performance Evaluation
As groundwork for directed graph generation, we prioritize strong unconditional generation, which we consider a key prerequisite for effective digraph modeling. At the same time, steering generation with graph-level properties is important for many downstream tasks. To illustrate the flexibility of our framework, we show that Directo can be extended to conditional generation without major architectural changes via classifier-free guidance (Ho and Salimans, 2022; Nisonoff et al., 2025). On the TPU Tiles dataset, we condition on the execution time of the computational graph, aiming at low execution times. Directo performs competitively with specialized autoregressive models like GraphRNN and LayerDAG (Appendix H.8), even surpassing the OneShotDAG variant of Li et al. (2025).
@@ 5.2 Ablation study
In the following, we present the results for ablations on the role of dual attention and the different positional encodings. We complement this with a scalability study and results on conditional generation. Complete results and further ablations on other model components can be found in Appendix H.
@@
width= Dataset Model Ratio $\downarrow$ V.U.N. $\uparrow$ ER-DAG RRWP - Double 4.9 $\pm$ 0.2 72.0 $\pm$ 9.8 RRWP - Dual 1.7 $\pm$ 0.1 94.0 $\pm$ 1.0 MagLap - Double 4.3 $\pm$ 0.7 80.0 $\pm$ 6.3 MagLap - Dual 1.3 $\pm$ 0.2 91.0 $\pm$ 2.5 SBM RRWP - Double 26.3 $\pm$ 2.5 0.0 $\pm$ 0.0 RRWP - Dual 1.8 $\pm$ 0.5 99.5 $\pm$ 1.0 MagLap - Double 14.2 $\pm$ 1.5 8.0 $\pm$ 7.5 MagLap - Dual 1.9 $\pm$ 0.3 77.0 $\pm$ 7.6
@@ 6 Related work
Initial methods include auto-regressive models (You et al., 2018; Liao et al., 2019), which grow the graph progressively by inserting nodes and edges, but require node ordering for feasibility. On the other hand, one-shot graph generative models such as VAEs (Kipf and Welling, 2016; Simonovsky and Komodakis, 2018; Jin et al., 2018; Ma and Zhang, 2021; Vignac and Frossard, 2022), GANs (De Cao and Kipf, 2018; Krawczuk et al., 2021; Martinkus et al., 2022), and normalizing flows (Kaushalya et al., 2019; Wehenkel and Louppe, 2021; Lippe and Gavves, 2021; Luo et al., 2021b) can directly generate full graphs in a single forward pass, thereby eliminating the need for predefined node ordering. Graph diffusion models also belong to this class, with the initial approaches consisting of adaptations of continuous state-space discrete-time diffusion frameworks (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song and Ermon, 2019) to the graph setting. These include works such as EDP-GNN (Niu et al., 2020), DGSM (Luo et al., 2021a), and GeoDiff (Xu et al., 2022). Later, the discrete state-space diffusion framework (Austin et al., 2021a) was adopted, with DiGress (Vignac et al., 2023a) and MCD (Haefeli et al., 2022) operating in discrete time. On the other hand, methods that operate in continuous time (Song et al., 2021; Campbell et al., 2022) were introduced, including GDSS (Jo et al., 2022) and GruM (Jo et al., 2024) in continuous state-spaces and DisCo (Xu et al., 2024) and Cometh (Siraudin et al., 2024) for discrete state-spaces. Recent approaches combine the sequential modeling of auto-regressive methods with the global denoising of diffusion (Zhao et al., 2024). Finally, DeFoG (Qin et al., 2025a) recently achieves state-of-the-art performance by leveraging discrete flow matching (Campbell et al., 2024; Gat et al., 2024).
@@ 7 Conclusion and future directions
Despite these promising results, limitations remain. First, scalability can be further improved in handling larger datasets and graph sizes (a direct result of the combinatorial complexity of digraphs). Incorporating scalable strategies, such as sparsity-aware attention (Qin et al., 2025b), hierarchical (Bergmeister et al., 2024; Jang et al., 2024), or latent diffusion (Yang et al., 2024), could help extend Directo at scale. Second, although our current setup includes classifier-free conditional generation, exploring alternative conditional strategies could further improve performance. Third, while Directo can learn structural properties such as acyclicity implicitly, enforcing such constraints (e.g., through PRODIGY (Sharma et al., 2024) or ConStruct (Madeira et al., 2024)) could further improve controllability and domain-specific validity. Finally, although the proposed graph transformer is used here for graph-to-graph generation, it can be readily adapted to discriminative tasks such as link prediction or node classification. Given its performance in the generative setting, it would be interesting to study how the architecture transfers to these alternative setups. Exploring these directions would enhance the applicability and impact of digraph generative models and our proposed architecture, particularly on new real-world scenarios. Appendix K includes a detailed discussion of limitations and future work.
Acknowledgements
This work was supported by the Swiss National Science Foundation (SNSF) under grant 10001445.
Reproducibility Statement
To ensure the reproducibility of our results, we have taken the following steps. The implementation of our method is provided in a GitHub repository, including all code and scripts necessary to reproduce the experiments and details on how to run it. In addition, we detail the proposed algorithm in Appendix D, while Appendix G provides full descriptions of the baselines, training setup, and hyperparameters employed. We also report the computational resources used and the runtime required for all experiments. Together, these materials are intended to enable researchers to fully replicate and build upon our work.
Ethics Statement
The objective of this work is to advance methodologies for the generation of directed graphs by introducing architectural mechanisms that explicitly model directionality and asymmetric relationships. Directed graphs are central to a wide range of applications, including causal inference, traffic modeling, and biological network analysis. Accordingly, improvements in directed graph generation have the potential to support progress in scientific research, decision-making systems, infrastructure design, or causal discovery.
The proposed framework strengthens the ability of generative models to handle directed graphs, and, at the moment, we do not identify any immediate societal risks associated with its deployment in its current form. Moreover, although the method is capable of generating structured and semantically meaningful graphs, the scale and complexity of these outputs remain limited, which may difficult direct applicability in high-impact domains such as clinical research or policy modeling at this stage.
Use of Large Language Models
Large Language Models (LLMs) were used as a writing assistant tool in the preparation of this manuscript. Specifically, they were employed to aid in writing, checking spelling, and editing for clarity. Nonetheless, all text produced with the assistance of LLMs was carefully reviewed, verified, and revised by the authors to ensure accuracy and appropriateness. The use of LLMs was limited to these support functions, and the authors take full responsibility for the final content of the paper.